OpenStudy (anonymous):

Someone please help i don't understand this at all! Suppose that gold is selling at $375/ounce. How many mmilligrams can you buy for one cent? (16 oz=1 lb; 1 lb=454 g)

OpenStudy (aaronq):

\(\dfrac{$375}{1\;oz.}=\dfrac{$0.01}{x\;oz.}\) then, \(\dfrac{1 oz.}{x \;oz.} = \dfrac{28.3495 g}{\gamma \;g}\) multiply the grams by 1000 to convert to milligrams.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."